# Solving Harder Reinforcement Learning Problems

### v1

@@ Abstract
Reinforcement Learning algorithms commonly sample multiple ($n>1$) solution attempts for each problem and reward them independently. This optimizes for $\mathrm{pass@1}$ performance and prioritizes the strength of isolated samples at the expense of the diversity and collective utility of sets of samples. Such algorithms under-utilize the sampling capacity, limiting exploration and eventual improvement on harder examples. As a fix, we propose Pass-at-$k$ Policy Optimization (PKPO), a transformation on the final rewards which leads to direct optimization of $\mathrm{pass@k}$ performance, thus optimizing for sets of samples that feature a large maximum reward when considered jointly. Our primary contribution is to derive novel low variance unbiased estimators for the $\mathrm{pass@k}$ and its gradient, in both the binary and continuous reward settings. We show that the optimization with these estimators reduces to standard reinforcement learning with (mini-batches of) rewards that have been jointly transformed by a function that is stable and efficient to compute.
@@
We validate our reward transformations on illustrative toy experiments, which reveal the variance reducing properties of our formulations. We also include real-world examples using the open-source model Gemma-2. We find that our transformation effectively optimizes for the target $k$. Furthermore, higher $k$ values enable solving more and harder problems, while annealing $k$ boosts both the $\mathrm{pass@1}$ and $\mathrm{pass@k}$. Crucially, for challenging task sets where conventional $\mathrm{pass@1}$ optimization stalls, our $\mathrm{pass@k}$ approach unblocks learning, likely due to better exploration by prioritizing joint utility over the utility of individual samples.
@@ 1 Introduction
Recent years have seen the rapid rise of large language models (LLMs) trained with internet scale pretraining data [26] with post training using both supervised fine-tuning [33] and reinforcement learning (RL) [1, 30, 3, 15]. The seminal paradigm of RL with human feedback [8] is limited by the human-derived data it is based on and the reward hacking issues that arise from the use of subjective signals more generally [2]. To unblock the path to artificial super intelligence (ASI), current work is focusing on grounded reward signals that are free of fine-grained human input as in code generation [28, 22, 13, 38, 16] and mathematics [20, 4, 11, 37].
The policy gradients family of RL methods [34] has proven effective in language model training [15]. To scale towards ASI, model training needs to tackle challenging RL task sets with no known solutions but for which correctness may be verified, as in formal mathematics environments [4]. In such settings, the RL training loop both updates model parameters and searches for solutions to problems at the continuously advancing frontier of model capabilities.
The specific search algorithm introduces a coupling between inference and model updates, which means that naively optimizing the expected single sample reward, or $\mathrm{pass@1}$ may be sub-optimal. While various inference-time search methods are possible [18, 19, 21, 35], simply taking multiple independent samples from the model has proven rather effective [24]. Our contribution is to allow coupling this simple search method with the model parameter updates by allowing the robust optimization the $\mathrm{pass@k}$ objective, which is the expected maximum reward over $k$ independent samples.
@@
The $\mathrm{pass@k}$ was championed by [10] who gave a popular unbiased estimator of the metric which we derive from a new perspective (to set up our gradient estimators) as Theorem 1 and generalize to continuous rewards in Theorem 3. Concurrently with our work [32] offered an elegant variance reduction method for the gradient of the $\mathrm{pass@k}$ which corresponds to the special case $n=k$ of our Equation 33. Interpreting that $\mathrm{pass@k}$ as a partial sort, [12] and [36] present elegant approximations that are rather general but less efficient in our setting. Others have provided variational approximations for handling the closely related Best-of-$N$ [9, 5] and other more general [6] inference-time algorithms. The contrasting idea of training a model to approximate the Best-of-$N$ prediction with a single sample was addressed by [27]. Our contribution can be interpreted as a generalization and variance reduction of [32]. For a general discussion of gradient estimation, variance reduction, and Monte Carlo, we recommend [23, 25].
@@
Our theme is constructing robust estimators of the $\mathrm{pass@k}$ and its gradient given $n\geq k$ samples by averaging (over all $\binom{n}{k}$ subsets of size $k$) simple estimators that are functions of $k$ samples. This is straightforward for binary rewards (Section 2), using the counting proof of Theorem 4. We generalize to continuous rewards using the key trick of assuming without loss of generality that the rewards are sorted, as in Section 3. Finally, we give baselining methods that require more involved derivations due to averaging over all subsets that do not include a given element (to retain unbiasedness) but which boil down to the same easy-to-apply results in Section 4, yielding our Pass-at-$k$ Policy Optimization (PKPO). We present toy experiments in Section 5.1 which demonstrate the variance reduction afforded by our estimators. Finally, Section 5.2 demonstrates that using our reward transformation solves more tasks and selectively optimizes $\mathrm{pass@k}$ through RL experiments on the LLM Gemma-2 [31], thus showcasing real-world impact.
@@
It is easy to adapt any policy gradient algorithm to use our results. Assume a vector $(g(x_{1}),g(x_{2}),\dots,g(x_{n}))^{\top}$ of per-sample rewards for a given task. For example, the $x_{i}$ could be model samples of source code addressing a specific task (which should be the same for all $n$ samples), and $g$ could provide a numeric score that measures how many tests the code passes, or an overall binary pass indicator, or some combination with additional stylistic or brevity terms, etc. Then in order to optimize the $\mathrm{pass@k}$ of Equation 1 (or the continuous analog $\mathrm{max}_{g}\mathrm{@k}$ of Equation 11) we need only jointly transform our vector of rewards using the sloo_minus_one function implemented on Line 56 of LABEL:code:listing, which maps $\mathbb{R}^{n}\mapsto\mathbb{R}^{n}$.
@@ 2.1 An Unbiased $\mathrm{pass@k}$ Estimator
where (6) holds because the sum on the r.h.s. of (5) is the number of subsets of size $k$ of the $(n-c)$ incorrect elements. Since averaging in this way retains unbiasedness, this completes the proof. ∎
@@ 2.2 An Unbiased $\mathrm{pass@k}$ Gradient Estimator
Given a mini-batch of $n$ i.i.d. samples $x_{1},x_{2},\dots,x_{n}$ from $p(x|\theta)$ with corresponding correctness labels $f_{i}\in\{0,1\}$, we want to optimize the $\mathrm{pass@k}$ w.r.t. the model parameters $\theta$. Letting $c=\sum_{i=1}^{n}f_{i}$ be the number of correct samples, we can use the gradient estimator
^$\widehat{\nabla}=\sum_{i=1}^{n}r_{i}\nabla_{\theta}\log p(x_{i}|\theta),~{}~{}~{}\textrm{where}~{}~{}~{}r_{i}=\begin{cases}\frac{k}{n}&\mathrm{if~{}}f_{i}=1\\ \frac{k}{n}\cdot\rho(n-1,c,k-1)&\mathrm{if~{}}f_{i}=0,\\ \end{cases}$ (8)
assigns more weight to correct samples, but gives some reward to incorrect samples to encourage exploration. The following well-known results will be used to show that (8) is unbiased.
@@ 3 Continuous Rewards
$\mathrm{max}_{g}\mathrm{@k}\equiv\mathbb{E}\left[\max\left(\left\{g(x_{i})\right\}_{i=1}^{k}\right)\right].$ (11)
3.1 An Unbiased $\mathrm{max}_{g}\mathrm{@k}$ Estimator
The following estimator for the $\mathrm{max}_{g}\mathrm{@k}$ is a direct analog of $\rho$: given $n\geq k$ i.i.d. samples, assuming w.l.o.g. that the rewards $g_{i}=g(x_{i})$ are sorted, so that $g_{1}\leq g_{2}\leq\dots\leq g_{n}$ the estimator is
@@ 3.1 An Unbiased $\mathrm{max}_{g}\mathrm{@k}$ Estimator
$\rho^{(g)}(n,c,k)$ is an unbiased estimator of the $\mathrm{max}_{g}\mathrm{@k}$.
@@
since $\mu_{i}$ is the number of subsets of $11,2,\dots,i-$ 1 of size $k-1$, which equals $\binom{i-1}{k-1}$. The sum starts at $k$ because all subsets of size $k$ include elements that are greater than or equal to $g_{k}$. ∎
@@
3.2 An Unbiased $\mathrm{max}_{g}\mathrm{@k}$ Gradient Estimator
@@
^$\widehat{\nabla}^{(g)}$ $=\sum_{i=1}^{n}{s_{i}}\nabla_{\theta}\log p(x_{i}|\theta),$ (18)
@@
$m_{ii}$ =$=\begin{cases}\binom{i-1}{k-1}&\mathrm{~{}if~{}}i\geq k-1\\ 0&\mathrm{otherwise,}\end{cases}$ (20)
@@
$m_{ij}$ =$=\begin{cases}\binom{j-2}{k-2}&\mathrm{~{}if~{}}(j>i)\wedge(j\geq k)\wedge(k\geq 2)\\ 0&\mathrm{otherwise.}\end{cases}$ (21)
@@
^$\widehat{\nabla}^{(g)}$ is an unbiased estimator of the gradient of the $\mathrm{max}_{g}\mathrm{@k}$:
$\mathbb{E}_{x_{1},x_{2},\dots,x_{n}\sim p(x|\theta)}\left[\widehat{\nabla}^{(g)}\right]=\nabla_{\theta}\mathrm{max}_{g}\mathrm{@k}.$ (22)
@@
The diagonals $m_{ii}$ are zero if $i<k$ since the largest element of any subset of size $k$ is at least $k$. If $i\geq k$ then we fix $i$ and are left with $i-1$ elements from which to choose $k-1$ which we can do $\binom{i-1}{k-1}$ ways in line with (20).
The $m_{ij}$ for $i\neq j$ are obtained by fixing $i$ and $j$ leaving $j-2$ elements $11,2,\dots,i-1,\dots,i+1,\dots,j-$ 1 from which to choose $k-2$ which we can do $\binom{j-2}{k-2}$ ways in line with (21). ∎
@@ 4.1 Leave-One-Out Baseline for the Simple Case
4.2 Leave-One-Out Baseline for $\mathrm{max}_{g}\mathrm{@k}$
Baselining the $s_{i}$ of (19) in this way introduces bias however, each $s_{i}$ depends on all $x_{1},\dots,x_{n}$. We instead apply LOO to the following form of $s_{i}$ that follows from Theorem 4 and the proof thereof:
@@ 4.2 Leave-One-Out Baseline for $\mathrm{max}_{g}\mathrm{@k}$
4.3 $\mathrm{max}_{g}\mathrm{@(k-1)}$ Leave-One-Out Baseline for $\mathrm{max}_{g}\mathrm{@k}$
The baseline $b_{i}^{(k)}$ is an average of the $\mathrm{max}_{g}\mathrm{@k}$ estimates over sets of size $k$. For a number of samples $n$ equal to $k$, there are no such subsets to construct the baseline. [32] recently overcame this issue for the specific case $n=k$ by using $\mathrm{max}_{g}\mathrm{@(k-1)}$ as the baseline statistic. We generalize their approach to $k<n$ and to averaging over all subsets by defining similarly to Equation 27
@@ 5.1 One-Dimensional Toy Example
(a) $g(x)$ and $\mathrm{max}_{g}\mathrm{@k}$.
(b) $g(x)$ and $\nabla_{\theta}\mathrm{max}_{g}\mathrm{@k}$.
Figure 1: The effect $k$ has on the optimal policy for a one-dimensional toy problem. The policy is normal with mean parameter $\theta$ and fixed standard deviation $0.1$. For the $\mathrm{max}_{g}\mathrm{@k}$ (left) the optimal $\theta$ corresponds to the horizontal position with maximum $\mathrm{max}_{g}\mathrm{@k}$. For the derivative (right), the optimal $\theta$ corresponds to the location of the zero crossing. For larger $k$ the optimal $\theta$ is more risk tolerant, allowing more samples to exceed one (with zero reward) in order to increase the chance of obtaining at least one sample close to, but less than one (with large reward).
@@
The optimal policy under the $\mathrm{max}_{g}\mathrm{@k}$ reward varies with $k$ (see Figure 1). The variance of our estimators is compared in Figure 8 where $s^{(\textrm{loo}-1)}$ is the strongest.
5.2 RL on Gemma-2

### v2

@@ Abstract
Reinforcement Learning algorithms commonly sample multiple ($n>1$) solution attempts for each problem and reward them independently. This optimizes for $\mathrm{pass@1}$ performance and prioritizes individual sample performance over the diversity and collective utility of a set of samples. Such algorithms under-utilize the sampling capacity, limiting exploration and eventual improvement on harder examples. As a fix, we propose Pass-at-$k$ Policy Optimization (PKPO), a multivariate transformation on batches of rewards which leads to direct optimization of $\mathrm{pass@k}$ performance, thus optimizing for sets of samples that feature a large maximum reward when considered jointly. Our primary contribution is to derive novel low variance unbiased estimators for the $\mathrm{pass@k}$ and its gradient, in both the binary and continuous reward settings. We show that optimizing with these estimators reduces to reinforcement learning with (batches of) rewards that have been jointly transformed by a function that is stable and efficient to compute.
@@
We validate our transformations on illustrative toy experiments, which reveal the variance reducing properties of our formulations. We also include real-world examples using the open-source models Gemma2 and Llama3.1. We find that our transformation effectively optimizes for the target $k$. Furthermore, higher $k$ values enable solving more and harder problems, while annealing $k$ boosts both the $\mathrm{pass@1}$ and $\mathrm{pass@k}$. Crucially, for challenging task sets where conventional $\mathrm{pass@1}$ optimization stalls, our $\mathrm{pass@k}$ approach unblocks learning, likely by improving exploration through the prioritization of joint utility over the utility of individual samples
@@ 1 Introduction
Recent years have seen the rapid rise of large language models (LLMs) trained with internet-scale pretraining data [33] with post training using both supervised fine-tuning [40] and reinforcement learning (RL) [1, 37, 3, 18]. The seminal paradigm of RL with human feedback [9] is limited by the human-derived data it is based on and the reward hacking issues that arise from the use of subjective signals more generally [2]. To enable progress toward superhuman capabilities, current work is focusing on grounded reward signals that are free of fine-grained human input as in code generation [35, 28, 15, 45, 19] and mathematics [25, 5, 13, 44].
The policy gradients family of RL methods [41] has proven effective in language model training [18]. To scale to new capabilities, model training needs to tackle challenging RL task sets with no known solutions but for which correctness may be verified, as in formal mathematics environments [5]. In such settings, the RL training loop both updates model parameters and searches for solutions to problems at the continuously advancing frontier of model capabilities.
The specific search algorithm introduces a coupling between inference and model updates, which means that naively optimizing the expected single sample reward, or $\mathrm{pass@1}$ may be sub-optimal. While various inference-time search methods are possible [22, 24, 27, 42], simply taking multiple independent samples from the model has proven rather effective [30]. Our contribution is to couple this simple search method with model parameter updates by enabling robust optimization of the $\mathrm{pass@k}$ objective, which is the expected maximum reward over $k$ independent samples.
(a) $g(x)$ and $\mathrm{max}_{\hskip-0.40251ptg\hskip-0.80504pt}\mathrm{@k}$.
(b) $g(x)$ and $\nabla_{\theta}\mathrm{max}_{\hskip-0.40251ptg\hskip-0.80504pt}\mathrm{@k}$.
Figure 1: The effect of $k$ on the optimal policy for a one-dimensional toy problem. The policy is normal with mean parameter $\theta$ and fixed standard deviation $0.1$. For the $\mathrm{max}_{\hskip-0.40251ptg\hskip-0.80504pt}\mathrm{@k}$ objective (left, defined in Equation 11) the optimal $\theta$ corresponds to the horizontal position with maximum $\mathrm{max}_{\hskip-0.40251ptg\hskip-0.80504pt}\mathrm{@k}$. For the derivative (right, the estimation of which is the focus of this paper), the optimal $\theta$ corresponds to the location of the zero crossing. For larger $k$ the optimal $\theta$ is more risk tolerant, allowing more samples to exceed one (getting zero reward) in order to increase the chance of obtaining at least one sample close to, but less than one (getting a large reward). See Section 5.1 for more details.
@@
The $\mathrm{pass@k}$ was championed by [11] who gave a popular unbiased estimator of the metric which we derive from a new perspective (to set up our gradient estimators) as Theorem 1, generalize to continuous rewards in Theorem 3, and provide additional characterisations in Corollaries 2 and 3. Concurrently with our work [39] offered an elegant variance reduction method for the gradient of the $\mathrm{pass@k}$ which corresponds to the special case $n=k$ of our Equation 33. Interpreting $\mathrm{pass@k}$ in terms of a partial sort, [14] and [43] present elegant approximations that are rather general but less efficient in our setting. Others have provided variational approximations for handling the closely related Best-of-NN [10, 6] and other more general [7] inference-time algorithms. The contrasting idea of training a model to approximate the Best-of-NN prediction with a single sample was addressed by [34]. Our contribution can be interpreted as a generalization and variance reduction of [39]. For a general discussion of gradient estimation, variance reduction, and Monte Carlo, we recommend [29, 31].
@@
Our theme is constructing robust estimators of the $\mathrm{pass@k}$ and its gradient given $n\geq k$ samples by averaging (over all $\binom{n}{k}$ subsets of size $k$) simple estimators that are functions of $k$ samples. This is straightforward for binary rewards (Section 2), using the counting proof of Theorem 4. We generalize to continuous rewards using the key trick of assuming without loss of generality that the rewards are sorted, as in Section 3. Finally, we give baselining methods that require more involved derivations due to averaging over all subsets that do not include a given element (to retain unbiasedness) but which boil down to the same easy-to-apply results in Section 4, yielding our Pass-at-$k$ Policy Optimization (PKPO). We present toy experiments in Section 5.1 which demonstrate the variance reduction afforded by our estimators. Finally, Section 5.2 demonstrates that using our reward transformation solves more tasks and selectively optimizes $\mathrm{pass@k}$ through RL experiments on Gemma2 [38] and Llama3.1 [16], showcasing real-world impact.
@@
It is easy to adapt any policy gradient algorithm to use our results. Assume a vector $(g(x_{1}),g(x_{2}),\dots,g(x_{n}))^{\top}$ of per-sample rewards for a given task. For example, the $x_{i}$ could be model samples of source code addressing a specific task (which should be the same for all $n$ samples), and $g$ could provide a numeric score that measures how many tests the code passes, or an overall binary pass indicator, or some combination with additional stylistic or brevity terms, etc. Then in order to optimize the $\mathrm{pass@k}$ of Equation 1 (or the continuous analog $\mathrm{max}_{\hskip-0.40974ptg\hskip-0.81949pt}\mathrm{@k}$ of Equation 11) we simply transform the vector of rewards using either the sloo or the sloo_minus_one function of LABEL:code:listing, which map $\mathbb{R}^{n}\mapsto\mathbb{R}^{n}$. [^1] While our experiments focus on sloo_minus_one, we recommend experimenting with both estimators.
@@ 2.1 An Unbiased $\mathrm{pass@k}$ Estimator
where (6) holds because the sum on the r.h.s. of (5) is the number of subsets of size $k$ of the (nc)(n-c) incorrect elements. Since averaging in this way retains unbiasedness, this completes the proof. ∎ We show in Corollary 2 that no such unbiased estimator exists for $n\leq k$, and in Corollary 3 that the asymptotic variance of this estimator decreases at a rate of $1/n$.
@@ 2.2 An Unbiased $\mathrm{pass@k}$ Gradient Estimator
Given a mini-batch of $n$ i.i.d. samples $x_{1},x_{2},\dots,x_{n}$ from $p(x|\theta)$ with corresponding correctness labels $f_{i}\in\{0,1\}$, we want to optimize the $\mathrm{pass@k}$ w.r.t. the model parameters $\theta$. Letting $c=\sum_{i=1}^{n}f_{i}$ be the number of correct samples, we will demonstrate unbiasedness of the estimator
[…]
that assigns more weight to correct samples, while also assigning some reward to incorrect samples to encourage exploration. The following well-known results will be used to show that (8) is unbiased.
@@ 3 Continuous Rewards
$\mathrm{max}_{\hskip-0.40974ptg\hskip-0.81949pt}\mathrm{@k}\equiv\mathbb{E}\left[\max\left(\left\{g(x_{i})\right\}_{i=1}^{k}\right)\right].$ (11)
3.1 An Unbiased $\mathrm{max}_{\hskip-0.40974ptg\hskip-0.81949pt}\mathrm{@k}$ Estimator
The following estimator for the $\mathrm{max}_{\hskip-0.40974ptg\hskip-0.81949pt}\mathrm{@k}$ is a direct analog of $\rho$: given $n\geq k$ i.i.d. samples, assuming w.l.o.g. that the rewards $g_{i}=g(x_{i})$ are sorted, so that $g_{1}\leq g_{2}\leq\dots\leq g_{n}$ the estimator is
@@
$\rho^{(g)}(n,c,k)$ is an unbiased estimator of the $\mathrm{max}_{\hskip-0.40974ptg\hskip-0.81949pt}\mathrm{@k}$.
@@
since $\mu_{i}$ is the number of subsets of $1,2,\dots,i-1$ of size k1k-1, which equals $\binom{i-1}{k-1}$. The sum starts at $k$ because all subsets of size $k$ include elements that are greater than or equal to $g_{k}$. ∎
@@
3.2 An Unbiased $\mathrm{max}_{\hskip-0.40974ptg\hskip-0.81949pt}\mathrm{@k}$ Gradient Estimator
@@
$\widehat{\nabla}^{(g)}$ $=\sum_{i=1}^{n}{s_{i}}\nabla_{\theta}\log p(x_{i}|\theta),$ (18)
@@
$m_{ii}$ $=\begin{cases}\binom{i-1}{k-1}&\mathrm{~if~}i\geq k-1\\ 0&\mathrm{otherwise,}\end{cases}$ (20)
@@
$m_{ij}$ $=\begin{cases}\binom{j-2}{k-2}&\mathrm{~if~}(j>i)\wedge(j\geq k)\wedge(k\geq 2)\\ 0&\mathrm{otherwise.}\end{cases}$ (21)
@@
$\widehat{\nabla}^{(g)}$ is an unbiased estimator of the gradient of the $\mathrm{max}_{\hskip-0.40974ptg\hskip-0.81949pt}\mathrm{@k}$:
$\mathbb{E}_{x_{1},x_{2},\dots,x_{n}\sim p(x|\theta)}\left[\widehat{\nabla}^{(g)}\right]=\nabla_{\theta}\mathrm{max}_{\hskip-0.40974ptg\hskip-0.81949pt}\mathrm{@k}.$ (22)
@@
The diagonals $m_{ii}$ are zero if $i<k$ since the largest element of any subset of size $k$ is at least $k$. If $i\geq k$ then we fix $i$ and are left with i1i-1 elements from which to choose k1k-1 which we can do $\binom{i-1}{k-1}$ ways in line with (20).
The $m_{ij}$ for $i\neq j$ are obtained by fixing $i$ and $j$ leaving j2j-2 elements $1,2,\dots,i-1,\dots,i+1,\dots,j-1$ from which to choose k2k-2 which we can do $\binom{j-2}{k-2}$ ways in line with (21). ∎
@@ 4.1 Leave-One-Out Baseline for the Simple Case
4.2 Leave-One-Out Baseline for $\mathrm{max}_{\hskip-0.40974ptg\hskip-0.81949pt}\mathrm{@k}$
Baselining the $s_{i}$ of (19) in this way introduces bias, however, as each $s_{i}$ depends on all $x_{1},\dots,x_{n}$. We instead apply LOO to the following form of $s_{i}$ that follows from Theorem 4 and the proof thereof:
@@
4.3 $\mathrm{max}_{\hskip-0.40974ptg\hskip-0.81949pt}\mathrm{@(k-1)}$ Leave-One-Out Baseline for $\mathrm{max}_{\hskip-0.40974ptg\hskip-0.81949pt}\mathrm{@k}$
The baseline $b_{i}^{(k)}$ is an average of the $\mathrm{max}_{\hskip-0.40974ptg\hskip-0.81949pt}\mathrm{@k}$ estimates over sets of size $k$. For a number of samples $n$ equal to $k$, there are no such subsets to construct the baseline. [39] recently overcame this issue for the specific case $n=k$ by using $\mathrm{max}_{\hskip-0.40974ptg\hskip-0.81949pt}\mathrm{@(k-1)}$ as the baseline statistic. We generalize their approach to $k<n$ and to averaging over all subsets by defining similarly to Equation 27
@@ 5.1 One-Dimensional Toy Example
The optimal policy under the $\mathrm{max}_{\hskip-0.40974ptg\hskip-0.81949pt}\mathrm{@k}$ reward varies with $k$ (see Figure 1). The variance of our estimators is compared in Figure 4 where $s^{(\textrm{loo}-1)}$ is the strongest.
Table 1: Results for Gemma2-9B on the MATH benchmark.
Gemma2-9B k_eval=1 k_eval=2 k_eval=4 k_eval=8 k_eval=16
k_opt=1 22.24 $\pm$ 0.50 25.35 $\pm$ 0.55 30.73 $\pm$ 0.59 37.08 $\pm$ 0.64 42.59 $\pm$ 0.68
k_opt=2 21.46 $\pm$ 0.51 28.61 $\pm$ 0.56 32.92 $\pm$ 0.61 39.59 $\pm$ 0.66 45.34 $\pm$ 0.70
k_opt=4 21.25 $\pm$ 0.53 27.15 $\pm$ 0.58 34.93 $\pm$ 0.63 41.71 $\pm$ 0.69 47.05 $\pm$ 0.74
k_opt=8 20.69 $\pm$ 0.56 26.78 $\pm$ 0.60 33.68 $\pm$ 0.66 42.62 $\pm$ 0.72 48.37 $\pm$ 0.77
[39] 19.48 $\pm$ 0.61 25.41 $\pm$ 0.67 31.17 $\pm$ 0.73 39.34 $\pm$ 0.79 44.82 $\pm$ 0.83
EntropyReg 20.85 $\pm$ 0.58 26.05 $\pm$ 0.64 32.48 $\pm$ 0.70 38.21 $\pm$ 0.76 43.95 $\pm$ 0.81
Table 2: Results for Llama3.1-8B on the MATH benchmark.
Llama3.1-8B k_eval=1 k_eval=2 k_eval=4 k_eval=8 k_eval=16
k_opt=1 51.15 $\pm$ 0.61 51.82 $\pm$ 0.64 53.69 $\pm$ 0.68 55.41 $\pm$ 0.72 56.83 $\pm$ 0.76
k_opt=2 49.72 $\pm$ 0.62 53.51 $\pm$ 0.66 55.45 $\pm$ 0.70 57.23 $\pm$ 0.74 58.71 $\pm$ 0.78
k_opt=4 49.18 $\pm$ 0.64 52.20 $\pm$ 0.68 57.83 $\pm$ 0.72 58.47 $\pm$ 0.77 59.28 $\pm$ 0.81
k_opt=8 48.63 $\pm$ 0.67 52.14 $\pm$ 0.71 56.28 $\pm$ 0.75 59.04 $\pm$ 0.80 61.88 $\pm$ 0.84
[39] 48.21 $\pm$ 0.70 50.93 $\pm$ 0.75 54.38 $\pm$ 0.80 57.11 $\pm$ 0.85 58.55 $\pm$ 0.90
EntropyReg 48.51 $\pm$ 0.68 51.95 $\pm$ 0.73 55.33 $\pm$ 0.78 56.95 $\pm$ 0.83 58.18 $\pm$ 0.88
Table 3: Results for Gemma2-9B on the Coding benchmark.
Gemma2-9B k_eval=1 k_eval=2 k_eval=4 k_eval=8 k_eval=16
k_opt=1 37.71 $\pm$ 0.60 42.03 $\pm$ 0.65 48.19 $\pm$ 0.69 55.07 $\pm$ 0.75 60.98 $\pm$ 0.79
k_opt=2 36.84 $\pm$ 0.61 46.56 $\pm$ 0.67 52.68 $\pm$ 0.72 59.73 $\pm$ 0.78 65.86 $\pm$ 0.84
k_opt=4 36.49 $\pm$ 0.63 44.95 $\pm$ 0.69 57.09 $\pm$ 0.76 63.64 $\pm$ 0.83 69.51 $\pm$ 0.88
k_opt=8 35.75 $\pm$ 0.67 44.41 $\pm$ 0.73 55.08 $\pm$ 0.80 65.56 $\pm$ 0.88 71.91 $\pm$ 0.94
[39] 34.36 $\pm$ 0.72 42.81 $\pm$ 0.78 52.36 $\pm$ 0.86 61.07 $\pm$ 0.95 66.41 $\pm$ 1.01
EntropyReg 35.91 $\pm$ 0.70 43.75 $\pm$ 0.76 53.28 $\pm$ 0.84 60.13 $\pm$ 0.93 65.29 $\pm$ 0.99
Table 4: Results for Llama3.1-8B on the Coding benchmark.
Llama3.1-8B k_eval=1 k_eval=2 k_eval=4 k_eval=8 k_eval=16
k_opt=1 67.38 $\pm$ 0.72 67.45 $\pm$ 0.76 69.22 $\pm$ 0.80 71.11 $\pm$ 0.84 72.84 $\pm$ 0.88
k_opt=2 64.91 $\pm$ 0.73 69.73 $\pm$ 0.78 72.03 $\pm$ 0.82 74.08 $\pm$ 0.87 75.89 $\pm$ 0.91
k_opt=4 64.25 $\pm$ 0.75 68.47 $\pm$ 0.80 74.67 $\pm$ 0.85 75.01 $\pm$ 0.90 77.75 $\pm$ 0.95
k_opt=8 63.57 $\pm$ 0.78 68.39 $\pm$ 0.83 72.84 $\pm$ 0.88 76.82 $\pm$ 0.94 79.33 $\pm$ 0.99
[39] 62.77 $\pm$ 0.82 66.86 $\pm$ 0.88 70.83 $\pm$ 0.94 73.95 $\pm$ 1.01 75.47 $\pm$ 1.06
EntropyReg 63.91 $\pm$ 0.80 67.85 $\pm$ 0.86 71.78 $\pm$ 0.92 72.99 $\pm$ 0.98 74.31 $\pm$ 1.04
5.2 RL on Open Source LLMs